OpenStudy (castiel):

integrate 1/sqrt(x^2-1) from 2 to 3?

OpenStudy (kirbykirby):

you have a form \(\sqrt{x^2-a^2}\) suggesting to use the trigonometric substitution \(x = \sec \theta\) \(dx = \sec \theta \tan \theta \, d\theta\) Ignoring the bounds for now, \[ \int \frac{1}{\sqrt{x^2-1}}dx=\int \frac{1}{\sqrt{\sec^2 \theta-1}}(\sec \theta \tan \theta \, d\theta)\\ \, \\ = \int\frac{\sec \theta \tan \theta}{\tan \theta}d\theta, \text{ since } \sec^2\theta -1 = \tan^2 \theta \\ \, \\ = \int \sec \theta d \theta \\ = \ln|\sec \theta +\tan \theta |+C\] Now into x-variables again, notice that \(x = \sec \theta\) and \(\sqrt{x^2-1}=\tan \theta\) from the simplification in your denominator, so \[=\ln\left|x+\sqrt{x^2-1}\right|+C \] You can drop the absolute value bars since you will be integrating from 2 to 3. With the bounds: \[ \large = \left [ \ln \left(x+\sqrt{x^2-1}\right) \right]_{2}^3 = \ln(3+\sqrt{3^2-1})-\ln(2+\sqrt{2^2-1}) \\ \large =\ln(3+2\sqrt{2})-\ln(2+\sqrt{3})\]